OpenStudy (anonymous):

suppose that an experiment is repeated four times. A certain event has probability 1/10 in a single repetition of the experiment. What is the probability that the event occurs every time in the 4 trials? What is the probability that it never occurs? What is the probability that it occurs at least once? What is the probability that it occurs in the fourth trial only?

OpenStudy (anonymous):

Ask this in the Mathematics section if it does not receive enough replies.

OpenStudy (matt101):

Probability of occurring 4 times = (1/10)^4 Probability of never occurring = (9/10)^4 Probability that it occurs at least once = 1 - (9/10)^4 Probability that it occurs in the fourth trial only = (9/10)^3 * 1/10 If you don't understand how I got any of those answers, ask :)

OpenStudy (anonymous):

these are simple Bernoulli trials..just read and use the formula..

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Bernoulli_trial